OpenStudy (anonymous):

The quadratic equation (p+q)x^2+px-5=0 has roots -5/2 and 1/3.Find the values of p and of q. @ganeshie8

ganeshie8 (ganeshie8):

setup two equation using sum of roots and product of roots ?

OpenStudy (anonymous):

THNX @ganeshie8

ganeshie8 (ganeshie8):

you're familiar with that stuff right ?

OpenStudy (anonymous):

yes :)

ganeshie8 (ganeshie8):

\[ax^2+bx+c=0\] sum of roots = -b/a product of roots = c/a

ganeshie8 (ganeshie8):

okay good :)

OpenStudy (anonymous):

:)